A 2.5-liter object (mass: 2 kg) is released from rest at a depth of 20 m below the surface of the ocean.

a) After it's released from rest, will the object rise to the surface or sink to the ocean floor? Explain.
b) Determine the object's acceleration after it's released.
c) Determine the speed of the object 3.5 seconds after it's released.
d) How far has the object traveled in that 3.5 seconds?
Answer #1

[a] ans

the onject rise to the surface to the ocean floor because bouancy force act on the object

[b] ans

now we find the acceleration

acceleration a=2*20=40 m/s^2

[c] ans

now we find the speed of the object

speed V=40*3.5=140 m/s

[d] ans

now we find the distance

distance S=0+1/2*40*3.5^2=245 m